TWO-PHOTON TRANSITIONS VIA OPTICAL FREQUENCY COMB TO FORM ULTRACOLD MOLECULES

\underline{SVETLANA MALINOVSKAYA}, WUFU SHI, Department of Physics and Engineering Physics, Stevens Institute of Technology, Hoboken, NJ 07030.

We show that a single femtosecond optical frequency comb may be used to induce two-photon transitions between molecular vibrational levels to form ultracold molecules, e.g., KRb. The phase across an individual pulse in the pulse train is sinusoidally modulated with a carefully chosen amplitude and modulation frequency. Piecewise adiabatic population transfer is fulfilled to the final state by each pulse in the applied pulse train providing a controlled population accumulation in the final state. Similar results are obtained using a standard optical frequency comb with zero carrier-envelop-offset frequency.